OpenStudy (anonymous):

find the length of the third side of a right triangle given the measurement of the other two sides. 7. 6 cm, 8 cm (1 point) 9 cm 7 cm 10 cm 11 cm

OpenStudy (anonymous):

do they tell you which sides they are?

OpenStudy (anonymous):

i mean are they the two shorter sides?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

10..

OpenStudy (anonymous):

all options are larger than 6 and 8 already..

OpenStudy (anonymous):



OpenStudy (anonymous):

ok then this will be a bit longer: so we know that either 8 will be our longest side, or that we need to find the hypotenuse: so if we assume that 8 is the longest side, then we can use pythagoras' thm to calculate the other short side: c^2=a^2+b^2 64= 36+b^2 b^2=28 b=sqrt(28) seeing as thats not one of our answers, we know now that 6 and 8 are the 2 shorter sides: so we set up pythagoras again: c^2=36+64 c^2=100 c=10 this is one our answers, so we know its right

OpenStudy (anonymous):

@cinar one of the options was 7, which meant that 8 could have been the biggest side, so we had to check and make sure that it wasn't

OpenStudy (anonymous):

I did not noticed that.. 6,8,10 special right triangle..